ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

17. Control Theory and Optimization

1991 MS Classification: not given Allende Alonso, Sira Maríıa, Universidad Habana, Cuba Approximation Approach for solving stochastic lot size problem Stochastic dynamic lot size problem, when the distribution of the demands is unknown, is studied. A deterministic approximative optimization problem is formulated by using the Cornish–Fisher expansion.